OpenStudy (anonymous):

What is the standard form of 2y=-4x-20?

OpenStudy (anonymous):

Ax+By=C -4x-2y=20

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Well i suppose they want the x-term to be positive, so you might be better off writing: 4x+2y=-20

OpenStudy (anonymous):

Or.... 2x + y = -10. Infinitely many versions of the same line in standard form.

OpenStudy (anonymous):

Not absolutely required, but I prefer to have the coefficients mutually prime integers if possible, although some lines won't allow it.

OpenStudy (anonymous):

...and as small as possible.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're very welcome.

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

In most cases yes, but that assumes the slope is a rational number.

OpenStudy (anonymous):

The first is preferred, but I don't think the second is wrong.

Directrix (directrix):

The coefficient of the x term should be positive as stated in the definition.

Directrix (directrix):

-4x-2y=20 is NOT the equation of a line in standard form.

OpenStudy (anonymous):

yes, i corrected myself up above

Directrix (directrix):

4x +2y=-20 is correct. A = 4, B = 2, and C = -20

OpenStudy (anonymous):

Okay, I'll put in the whole problem. Solve the following system of linear equations by addition: 6x+6y=-6 2y=-4x-20

Directrix (directrix):

What does this system have to do with the standard form question. I'm mixed up.

OpenStudy (anonymous):

I need to put 2y=-4x-20 into standard form to solve the system.

Directrix (directrix):

6x+6y=-6 2y=-4x-20 ------------ We don't need to worry about standard form. We need to eliminate a variable.

Directrix (directrix):

6x + 6y = -6 4x + 2y = -20 ---------------

OpenStudy (anonymous):

I know we do. Sorry. I thought I had to put it in standard form first. My teacher doesn't like computer program we are using.

Directrix (directrix):

We'll do that. The lines are in standard form above. I'll move them here. 6x + 6y = -6 4x + 2y = -20 ---------------

Directrix (directrix):

Are you supposed to use Addition/Subtraction or Substitution or some other technique to solve the system?

OpenStudy (anonymous):

Addition.

Directrix (directrix):

Look at the two y terms: 6y and 2y. What do you multiply 2y by to get 6y as a product?

Directrix (directrix):

Don't overthink. Two times ? = 6.

OpenStudy (anonymous):

3

Directrix (directrix):

Ok, we want the y terms to add to zero so we will multiply the second equation by -3.

OpenStudy (anonymous):

How did you know to multiply the second equation by -3?

Directrix (directrix):

I looked at the x terms and saw 6x and 4x. For them to add out, the first equation would have to be multiplied by 4 and the second by -6 to get coefficients of 24x and then -24x.

Directrix (directrix):

But, for the y terms to add out, only one multiplication is necessary. 4x + 2y = -20 becomes -3(4x + 2y) = -20 * (-3) -12x - 6y = 60 The y term of the second equation is ready to add out with the y term of the first equation.

Directrix (directrix):

6x+6y=-6 -12x - 6y = 60 --------------------- We are ready to add.

Directrix (directrix):

-6x + 0 = 54 -6x = 54 x = -9

Directrix (directrix):

Do you see that the point of multiplying by -3 is to eliminate one of the variables so that the other variable can be found?

OpenStudy (anonymous):

Yes. Weren't we supposed to multiply the first equation by something?

Directrix (directrix):

We could multiply it by 1. Any other number would change the y term's coefficent and cause the y-terms not to be eliminated. Now, if the system had been: 3x + 5y = 10 7x + 9y = 75 ------------------ then, two multiplications would be required.

Directrix (directrix):

We should solve for y.

OpenStudy (anonymous):

Okay, x=-9. So when multiplying equations, how can I tell when to multiply by 1?

Directrix (directrix):

4x + 2y = -20 4(-9) + 2y = -20 -36 + 2y = -20 2y = -20 + 36 2y = 16 y = 8 (-9,8)

Directrix (directrix):

Look at this system. What might be a plan to eliminate one variable? Your choice of variable. 3x + 5y = 10 7x + 9y = 75 ------------------

OpenStudy (anonymous):

I'm sorry, I don't know. I get confused when all the coefficients are positive. When something is negative, I can usually figure it out.

Directrix (directrix):

Which variable do you want to take out first. Choose one.

OpenStudy (anonymous):

The x variable.

Directrix (directrix):

3x + 5y = 10 7x + 9y = 75 ------------- Multiply the x term in the first equation by 7 and the x term in the second equation by 3. Then, write the results here.

OpenStudy (anonymous):

They are both 21.

Directrix (directrix):

3x + 5y = 10 7x + 9y = 75 ------------- Multiply each by 7 and stack up the result here.

OpenStudy (anonymous):

21x+35y=70 49x+63y=525

Directrix (directrix):

Good. If the terms were added as they are now, no variable would be eliminated. Do you agree?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

One of the original equations, or the ones that we have now that I've multiplied them by 7?

Directrix (directrix):

I should have said to multiply the first equation by 7 and the second equation by 3 to get the x terms the same. So, if you'll do that and post here. Sorry about the bad instructions. 3x + 5y = 10 7x + 9y = 75 ------------- 21x+35y=70-------> This has already been multiplied by 7 7x + 9y = 75 ------> Multiply by 3. ------------- 21x+35y=70 ? ------------

OpenStudy (anonymous):

21x+35y=70 21x+18y=225 _______________

Directrix (directrix):

Now, multiply your choice of one of the above equations by -1. As is, no variable will add out. I am going to post a summary of solving simultaneous equations here and also give the web address while you do the multiplying by -1.

OpenStudy (anonymous):

21x+18y=225 -21x-35y=-70 -------------- -17y=155 --- ---- -17 -17 y=9.12, which isn't right.

Directrix (directrix):

How do you know it is not right.

OpenStudy (anonymous):

Because 9.12 isn't a whole number. Variables are whole numbers. Yes, I could be wrong about that.

Directrix (directrix):

I asked because I made up the problem and did not know the answer myself. But, this brings up a good point. Variable do NOT have to be whole numbers or integers.

OpenStudy (anonymous):

Okay. I wasn't sure.

Directrix (directrix):

The answers here are (285/8 , -155/8) http://www.wolframalpha.com/input/?i=%283x+%2B+5y+%3D+10+and+%287x+%2B+9y+%3D+75%29

Directrix (directrix):

Some sets of simultaneous equations do not have a solution so be open about what a solution is.

OpenStudy (anonymous):

Okay. I know. Some of the problems in the program are inconsistent or dependent. Thank you.

Directrix (directrix):

I sense your frustration. Anybody who has studied math has experienced that. I'll add the web address for these guides. Purple Math is one of my favorite sites.

Directrix (directrix):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

Okay. Thank you.

Directrix (directrix):

Keep working.

OpenStudy (anonymous):

Okay.